Prev Next

SW&IT Meeting Minutes 2020-08-03

Attendees: David, Igor, Matthieu, Stephen, Trey, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200727
LR20200727 - Bec, Matthieu 40.0
LR20200727 - Holmberg, Carl 40.0
LR20200727 - Edmeades, David 24.0
LR20200727 - Sola, Igor 0.0
LR20200727 - Hooper, Stephen 80.0
LR20200727 - Bryant, Trey 0.0
LR20200727 - Zhang, Yang 42.0
LR20200727 - Zhang, Xianyu 32.0

News from management

  • 2020B schedule (draft) 2020B_V0.0.pdf: 2020B_V0.0.pdf
  • UITS visit at the summit to prepare VoIP

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • Laptop with OSSGUI for Mike G

Tip of the day

Team reports

YZ

  • Previous:
    • Postpone DMS LBC patch until after restart (so we can have telemetry flowing again and do some tests).
    • Some TCS discussions (astyle, new release around SSD restart, etc.).
    • TMS meeting and Ciddor correction verification with Trenton (lots of software work in the next 2-3 weeks, due to limited window in SSD for TMS activities).
    • AO-Cravat (OSCO data analysis distributed to wide group of people in LBTO, some discussions, some research and studies on enhancing our Strehl code, meeting etc.).
    • Wiki work for both TMS and AO-Cravat.
    • Finally did some safety training!
  • Next:
    • Mostly will focus on TMS software for the next week:
      • Continue Ciddor correction verification with Trenton.
      • Develop all proposed software components.
    • Some AO-Cravat work if have additional time.

DE

  • Last week short
  • Reopened Syno ticket, following up today
  • Wednesday Synology work in office, no complaints about timing yet
    • Visit request sent to CV
  • Safety training
  • Test GUI VM set up for MB
  • Jim has "lost his Anyconnect icon", working on directing him to the start menu....
  • Backup jobs hosted on Windows machines needed a kick to start back up after the power outage. They are all going now.
  • Cisco training ongoing

MB

  • Last week
    • OC, Management
    • PE's
    • Recruitment, schedule screening
    • TCS make astyle
    • resurrect python-iqe
    • resurrect demo gui for b/w tests
    • PEPSI new servers follow-up
    • sciarch meeting
  • This week
    • PE's complete
    • Recruitment screening
    • remote obs tests with JP
    • UITS wired security
    • fiber follow-up
    • safety training
    • DMS
      • websockets, AdSec, etc.
      • LBC stiched png preview
      • max_bytes for ovms

SH

  • Fully transitioned employees to new VPN * updated firmware on both vpn.tucsonn and vpn.lbto
  • Leroy recabled mountain VPN for me * recabling seems to have fixed failover issues * updated firmware on active/passive firewalls * about five seconds of downtime as core-b was unplugged and plugged back into a different port!
  • wired network security meeting, plans

next:

  • start migrating affiliates to new VPN device, and statify in Github
  • updates to obs/robs, etc. vis-a-vis John's scripts
  • emacs issue on shell
  • mountain firewall changes for PLC on 10.144.0.0/24 access
  • wired security meetings and planning

TB

  • Last week:
    • Able to graph composite (least noisy) data on last plot w/ correct colors
      • As scatter only - need to fix
    • Made small change in how the time period is formatted -> "hh:mm" string instead of integer
    • Made default plot window larger - more easily readable now
    • Fixed formatting so that titles and axis labels don't overlap
    • Attempted to resolve UTC time issue
  • This week:
    • Still to read about sensors/equipment
    • Fix for colormap scatter/plot issue for composite data
    • Next/prev functionality now that I have better understanding of its purpose
    • Make plot and prompt more interactive?

IS

Last week:
  • LUCI
    • Investigation of MCU WebIO for LUCI1. Also created a basic java test example, using the same logic as the LUCI code does to test the new WebIO and it seems to work. The new WebIOs can export the configuration so I will generate a configuration file for each of the 6 WebIOs we have, even though we have only two new ones, but having the configuration files already ready, will be a help. Also I still need to play with it a little bit. Like load it from a DCHP default configuration and then load the configuration, then reset so it can got back to DHCP and load a new configuration if necessary. Document that.
    • B4 vacation: "After uploading new masks in a cabinet exchange, somehow the old masks are still reported. I managed to replicate the problem with a lot of back and forth, trying to replicate the sequence as in the logs. This issue has been there since the beginning. I started developing a solution to it, trying to make it as small as possible." ... still working on this.

  • PIT
    • Created a new script to report on proposals by instrument rather than partner. Specifically PEPSI and LBTI. While at it fixed a couple of bugs/features on the xml2list generation, concerning targets

This week:
  • Safety Training
  • LUCI
    • Finish mask exchange issue.
    • Finish configuration of LUCI WebIO * RPM
    • I'd like to start moving the ball on this.
  • PIT
    • Work with Michelle on new PIT release
    • Work on ways to apply an RPM to it, even though the PIT might not be the best example application but it should be very similar to the OT.

XZ

Topic revision: r10 - 10 Aug 2020, XianyuZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback